# Axis-Aligned Document Dewarping

### v1

@@ 4.3 Implementation Details
Our model is implemented using the PyTorch framework [28]. The dewarping model is trained on two datasets: Doc3D [5] (88k images) and UVDoc [36] (20k images). The input images are resized to 712$\times$488 pixels, while the output 2D grid is set to 45$\times$31. We optimize the model using the AdamW optimizer [21] with a batch size of 72. The initial learning rate is $1\times 10^{-5}$ and is linearly decayed to 0 during training. The network is trained for 30 epochs on two NVIDIA 4090 GPUs, which takes approximately 15 hours. During inference (see Sec. 3.3), a single round of preprocessing is applied for the DocUNet and UVDoc benchmarks, whereas for the DIR300 benchmark—with its smaller target objects—preprocessing is performed twice.
@@ 4.4 Experimental Results
We compare our method against existing approaches both quantitatively and qualitatively on the aforementioned benchmarks.
@@
Table 1: Quantitative evaluation on the DocUNet benchmark [23]. We compare our method with existing methods in terms of Multi-Scale Structural Similarity (MS), Local Distortion (LD), Aligned Distortion (AD), Axis-Aligned Distortion (AAD), Character Error Rate (CER) and Edit Distance (ED), “↑” indicates the higher the better and “↓” means the opposite, Bold font denotes the best result, and underline indicates second-best. The “Improvements” row shows the relative gain of our metric over the previous best.
@@
DewarpNet[5] 0.474 8.362 0.398 0.164 824.5 0.225
DDCP[40] 0.472 8.982 0.428 0.159 1124.7 0.280
DocTr[7] 0.509 7.773 0.369 0.151 708.6 0.185
PaperEdge[24] 0.473 7.967 0.368 0.121 731.4 0.183
DocGeoNet[8] 0.504 7.717 0.381 0.158 695.0 0.187
FAM[16] 0.497 8.416 0.377 0.151 688.3 0.176
Layout[17] 0.525 6.706 0.300 0.121 689.8 0.180
UVDoc[36] 0.545 6.827 0.316 0.125 754.2 0.193
Ours 0.543 6.249 0.278 0.099 603.1 0.150
Improvements -0.4% 6.8% 7.3% 18.2% 12.4% 14.8%
[…]
@@
Figure 7: Qualitative evaluation on the DocUNet benchmark. The figure shows the dewarped results along with the corresponding heatmap overlays of the AAD component metric. The horizontal ($\leftrightarrow$) and vertical ($\updownarrow$) arrows indicate the directions for visualizing the axis-aligned error of the dewarped feature lines.
In Tab. 1, we compare our method with previous approaches on the DocUNet benchmark. Our approach demonstrates significant improvements across most metrics and achieves performance comparable to UVDoc [36] on the MS metric. Notably, our method delivers an 18.2% improvement on the proposed AAD metric. Furthermore, we performed a qualitative analysis of global images (see Fig. 7) by selecting several test cases rich in horizontal and vertical features for dewarping comparison. To highlight the distortion areas more clearly, we overlay the dewarped results from different methods with heatmaps generated from the horizontal and vertical components of the proposed AAD metric (an example is provided in Fig. 4). For our method, the resulting heatmaps appear fainter on documents with feature lines in varying orientations, indicating a reduced axis alignment error. As expected from the dewarping results in Fig. 1, the feature lines exhibit improved axis alignment in the dewarped images.
Table 2: Quantitative evaluation on the DIR300 benchmark[8]. The evaluation protocol is identical to that described in Tab. 1.
Method MS↑ LD↓ AD↓ AAD↓ ED↓ CER↓
DewarpNet[5] 0.492 13.944 0.332 0.147 1076.8 0.336
DDCP[40] 0.552 11.02 0.357 0.132 739.3 0.257
DocTr[7] 0.616 7.189 0.255 0.107 698.4 0.211
PaperEdge[24] 0.584 7.968 0.255 0.091 447.4 0.177
DocGeoNet[8] 0.638 6.403 0.241 0.106 672.5 0.206
FAM[16] 0.607 7.680 0.243 0.108 639.4 0.198
Layout[17] 0.652 5.702 0.195 0.087 495.4 0.173
UVDoc[36] 0.621 7.730 0.219 0.101 614.0 0.237
Ours 0.702 4.261 0.131 0.057 405.8 0.132
Improvements 7.7% 25.3% 32.8% 34.5% 9.3% 23.7%
[…]
Figure 8: Qualitative evaluations on the DIR300 benchmark. We compare our method with existing methods as described in Fig. 7.
In Tab. 2, we evaluate our method on the DIR300 benchmark, which comprises 300 images—the largest among the three benchmarks. Our approach achieves notable improvements across all metrics, with the AAD metric improvement to 34.5%. Furthermore, the dewarped results significantly boost OCR performance, with the CER reaching 23.7%. These improvements are qualitatively demonstrated in Fig. 8. Our method substantially reduces the distortion of document feature lines, resulting in fainter heatmap colors and yielding more complete and straighter contours.
Table 3: Quantitative evaluation on the UVDoc benchmark [36]. The evaluation protocol is identical to that described in Tab. 1.
Method MS↑ LD↓ AD↓ AAD↓ ED↓ CER↓
DewarpNet[5] 0.589 5.666 0.193 0.085 297.8 0.117
DDCP[40] 0.585 11.582 0.290 0.109 251.8 0.094
DocTr[7] 0.697 8.146 0.161 0.075 110.8 0.034
DocGeoNet[8] 0.706 8.350 0.168 0.072 123.1 0.047
RDGR[12] 0.610 6.423 0.281 0.096 122.45 0.051
UVDoc[36] 0.785 6.810 0.119 0.052 122.7 0.046
Ours 0.800 5.834 0.098 0.040 104.3 0.036
Improvements 1.9% -3.0% 17.6% 23.1% 5.9% -5.9%
[…]
Figure 9: Qualitative comparisons on the UVDoc benchmark. We compare our method with existing methods as described in Fig. 7.
In Tab. 3, we present the quantitative results on the UVDoc benchmark. Our method outperforms competing approaches across all metrics except for LD and CER, where our results are very close to the best values. Notably, the AAD metric improves by 23.1%. In the qualitative comparison shown in Fig. 9, dewarped results in both horizontal and vertical directions are displayed. These comparisons reveal that our method achieves excellent axis-aligned dewarping performance even in regions without textual features, highlighting the superiority of our approach.
4.5 Ablation Study
We conducted ablation experiments to evaluate the impact of two key components: the axis-aligned geometric constraint used during training (Sec. 3.1) and the axis alignment preprocessing applied during inference(Sec. 3.3).
As summarized in Tab. 4, both components independently enhance performance across three benchmark datasets. Specifically, for the DocUNet and UVDoc benchmarks—where a large proportion of targets are present—the axis-aligned geometric constraint yields greater performance improvements than the axis alignment preprocessing. Conversely, in the DIR300 benchmark, which contains a smaller proportion of targets, the benefits of the preprocessing strategy are more pronounced. Importantly, the combination of both modules produces the best overall performance, demonstrating their complementary advantages in different scenarios.
Table 4: Ablation study on various benchmarks. We evaluate the impact of the axis-aligned geometric constraint loss (AL) and the axis alignment preprocessing strategy (AP) module. The symbols ✗ and ✓ indicate that the corresponding module is disabled or enabled, respectively. Bold font denotes the best performance, while underline values indicate the second-best scores.
[…]
DocUNet[23] 0.538 6.897 0.322 0.124 638.1 0.167
0.549 6.566 0.291 0.105 718.0 0.183
0.536 6.767 0.308 0.120 628.5 0.165
0.543 6.249 0.278 0.099 603.1 0.150
DIR300[8] 0.615 7.691 0.215 0.100 633.0 0.243
0.629 7.880 0.176 0.076 498.9 0.206
0.694 4.303 0.152 0.069 541.6 0.153
0.702 4.261 0.131 0.057 405.8 0.132
UVDoc[36] 0.794 6.356 0.118 0.052 120.4 0.041
0.785 5.903 0.116 0.044 105.2 0.035
0.790 6.024 0.111 0.049 118.2 0.043
0.800 5.834 0.098 0.040 104.3 0.036
Fig. 10 provides several examples from our ablation experiments. For targets with a smaller proportion (middle row), the preprocessing strategy improves target integrity, while the addition of the axis-aligned geometric constraint further enhances the dewarping performance. For targets with a larger proportion (first and last rows), the performance gains from the axis-aligned geometric constraint are more prominent. Overall, the optimal dewarped results are achieved by applying both components simultaneously, which aligns with the improvements observed in the qualitative metrics in Tab. 4.
[…]
Figure 10: Dewarped results under different ablation settings. From left to right, each row displays the distorted input image followed by the dewarped outputs generated by different ablation settings. The heatmap value is the horizontal component of the AAD metric.
5 Conclusion
This paper proposes a document dewarping method that exploits the inherent axis-aligned properties of planar documents’ discrete grid geometry. By incorporating an axis-aligned geometric constraint during network training and applying an axis alignment preprocessing step during inference, our method outperforms state-of-the-art methods across multiple benchmarks. Experimental results demonstrate that these components are both independently effective and complementary. We also introduce a novel metric, AAD, which offers a clear geometric interpretation, aligns well with human visual perception, and exhibits enhanced robustness. Unlike previous studies that primarily focus on network architecture or dataset contributions, our approach leverages the intrinsic geometry of documents to improve dewarping performance. Future work will further explore these geometric properties and extend their application to other image dewarping tasks.

### v2

@@ 4.3 Implementation Details
[…]
@@ 4.4 Experimental Results
We quantitatively and qualitatively evaluate our method against others on the DocUNet and DIR300 benchmarks.
Table 1 summarizes the quantitative evaluation. On the DocUNet benchmark, our approach demonstrates significant improvements across most metrics, achieving performance comparable to UVDoc (Verhoeven et al. 2023) on the MS metric and delivering a notable 18.2% improvement on the proposed AAD metric. The improvements are even more pronounced on the larger DIR300 benchmark, where our method achieves gains across all metrics, including a 34.5% improvement in AAD and a significant boost to OCR performance, with the CER reaching 23.7%.
Qualitative results for both benchmarks are shown in Figure 6. To highlight distortion areas more clearly, we overlay the dewarped results with heatmaps generated from the horizontal and vertical components of our AAD metric (an example is provided in Figure 4). For our method, the resulting heatmaps appear consistently fainter on documents from both datasets, indicating substantially reduced axis-alignment error. As expected from the dewarping results in Figure 1(a), this yields more axis-aligned feature lines in the final dewarped images.
Figure 6: Qualitative evaluation. The figure shows the dewarped results along with the corresponding heatmap overlays of the AAD component metric. The horizontal ($\leftrightarrow$) and vertical ($\updownarrow$) arrows indicate the directions for visualizing the axis-aligned error of the dewarped feature lines.
Figure 7: Qualitative results of our dewarping ablation study. From left to right: the distorted input and dewarped outputs from different settings. The overlaid heatmaps visualize the horizontal component of the AAD metric.
4.5 Ablation Study
We conducted ablation experiments to evaluate the impact of two key components: the axis-aligned geometric constraint used during training (Section 3.1) and the axis alignment preprocessing applied during inference(Section 3.3).
As summarized in Table 1, both components independently enhance performance across two benchmark datasets. Specifically, for the DocUNet benchmark—where a large proportion of targets are present—the axis-aligned geometric constraint yields greater performance improvements than the axis alignment preprocessing. Conversely, in the DIR300 benchmark, which contains a smaller proportion of targets, the benefits of the preprocessing strategy are more pronounced. Importantly, the combination of both modules produces the best overall performance, demonstrating their complementary advantages in different scenarios.
Figure 7 provides several examples from our ablation experiments. For targets with a smaller proportion (middle row), the preprocessing strategy improves target integrity, while the addition of the axis-aligned geometric constraint further enhances the dewarping performance. For targets with a larger proportion (first and last rows), the performance gains from the axis-aligned geometric constraint are more prominent. Overall, the optimal dewarped results are achieved by applying both components simultaneously, which aligns with the improvements observed in the qualitative metrics in Table 1.
5 Conclusion
This paper proposes a document dewarping method that exploits the inherent axis-aligned property of planar documents’ underlying grid geometry. To this end, we introduce two key components: an axis-aligned geometric constraint for training and an alignment preprocessing strategy for inference. Experimental results confirm that these components are both independently effective and complementary. As a result, our full model outperforms state-of-the-art methods, demonstrating both significant quantitative improvements and clear qualitative advantages. To apply our geometric insight to the evaluation process itself, we introduce a novel evaluation metric AAD, which provides a clear geometric interpretation, aligns well with human visual perception, and exhibits enhanced robustness. Unlike previous work focused on network architectures or datasets, our approach leverages the intrinsic geometry of documents to improve dewarping performance. This principle-driven strategy demonstrates the significant value of exploiting such properties. Future work will involve extending this principle to other image rectification tasks.
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62495092), the Natural Science Basic Research Plan of Shaanxi Province, China (No. 2025SYS-SYSZD-023), and the China Scholarship Council (No. 202406280284).
Supplementary Material
OVERVIEW
In this supplementary material, we provide the following additional information:
  1. (1)
    Details of the Inference Process.
  2. (2)
    Details of the evaluation metrics.
  3. (3)
    Robustness analysis of the AAD Metric.
  4. (4)
    Comparison on UVDoc Benchmark.
  5. (5)
    Additional visual comparison of the AD and AAD metrics.
  6. (6)
    Visual comparison of OCR results on dewarped document images.
  7. (7)
    Additional experimental results.
Appendix A Details of the Inference Process
During inference, we first rotate and crop the document region using the axis alignment preprocessing method described in the main text. To ensure that the entire document area is preserved, we employ a region expansion strategy instead of merely clipping the edges based on the minimum rotation matrix boundary.
The expansion ratio is adjusted for each benchmark: for the UVDoc benchmark (Verhoeven, Magne, and Sorkine-Hornung 2023), it is set to 0.03; for the DIR300 benchmark (Feng et al. 2022), an initial expansion of 0.01 is followed by a secondary expansion of 0.03; and for the DocUNet benchmark (Ma et al. 2018), the ratio is set to 0.06.
Appendix B Details of the Evaluation Metrics
The multi-scale structural similarity (MS-SSIM) metric extends SSIM by evaluating image similarity at multiple scales and combining these measurements through a weighted average. We adopt the original weights from (Wang, Simoncelli, and Bovik 2003): [0.0448, 0.2856, 0.3001, 0.2363, 0.1333].
Local Distortion (LD) is computed using dense SIFT flow mapping (Liu, Yuen, and Torralba 2010) from the ground truth image to the dewarped image, and is defined as the average L2 distance between corresponding pixels (You et al. 2017). This metric effectively measures the average local distortion of the dewarped image relative to the ground truth. Aligned Distortion (AD) is a more robust variant introduced in (Ma et al. 2022). Unlike LD, AD mitigates errors caused by global image translation and scaling by extracting the optimal affine transformation from the SIFT flow. It further weights the error according to the image gradient magnitude, emphasizing regions with significant features such as text or edges. In addition, the AAD metric proposed in the main text offers enhanced geometric interpretability and robustness compared to AD. We compute SIFT flow using the official implementation in MATLAB 2024b. Prior to computing these metrics, all dewarped and reference images are resized to a total area of 598,400 pixels, as recommended in (Ma et al. 2018).
Beyond image similarity metrics, we also evaluate OCR performance using character error rate (CER) and edit distance (ED) (Navarro 2001). Edit distance quantifies the difference between two character sequences, while CER is defined as the ratio of the edit distance to the total number of characters in the reference text, i.e., (s + i + d)/N, where s, i, and d denote the number of substitutions, insertions, and deletions, respectively, and N is the total number of characters in the reference text.
Appendix C Robustness Analysis of the AAD Metric
To validate the robustness of the proposed AAD metric, we applied geometric and illumination disturbances to the dewarped images shown in Figure S1, generating 100 warped images for each setting. To create a ground truth for evaluation, we calculate the ’true’ metric values using the known geometric disturbances as the ground truth flow. We then compute the ’reference’ metric values using the flow field estimated by the SIFT-flow (Liu, Yuen, and Torralba 2010) method. The correlation between these true and reference values indicates the metric’s robustness.
As shown in Figure S2(a), the AAD metric demonstrated the highest linear correlation ($R^{\text{2}}$) across all sets, indicating a strong linear relationship with the ground truth even under minimal geometric disturbances(i.e., in regions with lower metric values, which are critical for comparing dewarped results that closely approximate the true values). In contrast, the AD metric exhibited a nonlinear response in these regions.
In Figure S2(b), we evaluated the normalized standard deviation under three different illumination conditions for each image in Figure S1. The results revealed that the standard deviation for AAD (0.0230) was smaller than that for AD (0.0252), further confirming its enhanced robustness.
Figure S1: Geometric and illumination disturbances applied to the dewarped image. The leftmost column shows the input distorted image, alongside its corresponding dewarped and ground truth versions. On the right, geometric and illumination disturbances are added to the dewarped image, and the intensity increases along with the corresponding arrow direction. Set1 includes only geometric disturbances, Set2 adds color interference to Set1, and Set3 adds a random shadow to Set2.
Figure S2: Robustness analysis of each metric. (a) The x-axis represents the ground truth metric computed from the ground truth optical flow, while the y-axis represents the reference metric derived from the SIFT-flow method; ideally, they should exhibit a high linear correlation. (b) The normalized standard deviation of the metrics under three different environmental conditions, lower values indicate greater robustness against environmental disturbances.
Appendix D Comparison on UVDoc benchmark
This section provides a detailed analysis of our experimental results on the UVDoc benchmark (Verhoeven et al. 2023). UVDoc is a synthetic dataset generated via a pseudo-photorealistic rendering pipeline. This benchmark consists of 50 images, including matching pairs with and without illumination. Following the UVDoc protocol, 20 rich-text images were selected for OCR evaluation.
In Table S1, we present the quantitative results on the UVDoc benchmark. Our method outperforms competing approaches across all metrics except for LD and CER, where our results are very close to the best values. Notably, the AAD metric improves by 23.1%. The qualitative comparison in Figure S3 displays dewarped results for both horizontal and vertical directions. These comparisons reveal that our method achieves excellent axis-aligned dewarping performance even in regions without textual features, highlighting the superiority of our approach.
The corresponding ablation results in Table S1 are consistent with the other two benchmarks in the main text. The two proposed modules are individually effective, and they can be combined to effectively enhance performance.
Appendix E Additional Visual Comparison of the AD and AAD metrics
In Figure S4, we have provided additional case comparisons between the AAD and the AD metric. Each row contrasts the dewarped results obtained using UVDoc (Verhoeven et al. 2023) with those produced by our method, accompanied by corresponding heatmap visualizations for each metric. Notably, the differences observed in Figure S4(b) are primarily due to the axis alignment of text lines. These differences are distinctly captured by the AAD heatmap in Figure S4(d), reinforcing that the AAD metric more closely aligns with human perception of dewarping quality.
In contrast, the AD heatmap in Figure S4(c) lacks such discriminative features. Its color variations do not convey clear, meaningful distinctions, and the numerical values indicate that AD fails to effectively differentiate subtle differences. For example, the AD metric assigns a lower (better) score to the result on the left than the one on the right, a conclusion that contradicts visual assessment.
Our proposed AAD metric, however, successfully captures these subtle differences. When dewarping results deviate only slightly from the ground truth, the AD metric struggles to capture these nuances and does not maintain a consistent linear relationship with the true values—a limitation that could hinder future method improvements. In contrast, the AAD metric is more robust, interpretable, and closely aligned with human perception, making it an excellent evaluation metric for dewarping tasks.
Appendix F Visual Comparison of OCR Results on Dewarped Document Images
Figure S5 illustrates several cases demonstrating how dewarping results from different methods affect OCR performance. In each instance, we overlay the detected character bounding boxes on the dewarped images. Our method reliably recognizes text in nearly all regions, outperforming alternative approaches. Moreover, the minimal overlap between adjacent text boxes underscores the superior correction quality, yielding outputs that closely resemble the ground truth. Below each image is the computed OCR metric value, which shows a significant reduction in errors compared to the distorted input. These results confirm that improved dewarping directly contributes to lower OCR error rates, bringing the output closer to the ideal ground truth.
Appendix G Additional Experimental Results
Owing to space limitations in the main text, we include additional dewarping results in this supplementary material. Figure S6Figure S7, and Figure S8 present sample outputs on the DocUNet (Ma et al. 2018), DIR300 (Feng et al. 2022), and UVDoc (Verhoeven et al. 2023) benchmarks, respectively. These examples span a diverse range of document layouts. To facilitate a detailed assessment of the correction accuracy, we overlay the horizontal component heatmap of our AAD metric onto the dewarped images. As shown in the figures, our method consistently yields less distorted results with fainter heatmap intensities—indicating lower correction errors—compared to other approaches. These additional results further validate the robustness and superiority of our proposed method.
@@ Appendix G Additional Experimental Results
Table S1: Quantitative evaluation and ablation study on the UVDoc benchmark. We compare our method with previous methods and ablate our components: Axis-aligned geometric constraint Loss (AL) and Axis alignment Preprocessing (AP). “↑” indicates higher is better, “↓” means lower is better. Bold denotes the best result, and underline indicates second-best. The “Improvements” row shows the relative gain of our full model over the previous state-of-the-art.
@@
DewarpNet(Das et al. 2019) 0.589 5.666 0.193 0.085 297.8 0.117
DDCP(Xie et al. 2021) 0.585 11.582 0.290 0.109 251.8 0.094
DocTr(Feng et al. 2021) 0.697 8.146 0.161 0.075 110.8 0.034
DocGeoNet(Feng et al. 2022) 0.706 8.350 0.168 0.072 123.1 0.047
RDGR(Jiang et al. 2022) 0.610 6.423 0.281 0.096 122.45 0.051
UVDoc(Verhoeven et al. 2023) 0.785 6.810 0.119 0.052 122.7 0.046
  Ours (Baseline, w/o AL, AP) 0.794 6.356 0.118 0.052 120.4 0.041
  Ours (+ AL only) 0.785 5.903 0.116 0.044 105.2 0.035
  Ours (+ AP only) 0.790 6.024 0.111 0.049 118.2 0.043
Ours (Full, AL+AP) 0.800 5.834 0.098 0.040 104.3 0.036
Improvements +1.9% -3.0% +17.6% +23.1% +5.9% -5.9%
@@
Figure S3: Qualitative comparisons on the UVDoc benchmark. The figure shows the dewarped results along with the corresponding heatmap overlays of the AAD component metric. The horizontal ($\leftrightarrow$) and vertical ($\updownarrow$) arrows indicate the directions for visualizing the axis-aligned error of the dewarped feature lines.
Figure S4: Visualization of the AD and AAD metrics. (a) Ground truth image. (b) Comparison of dewarping results: the left image is produced by UVDoc (Verhoeven et al. 2023) and the right image is generated by our method. (c) Heatmap of the AD metric computed for the dewarped images relative to the ground truth. (d) Heatmap corresponding to the AAD metric proposed in the main text.
Figure S5: OCR results for dewarped documents produced by various methods. Each row overlays the detected text regions on the dewarped image derived from the distorted input (first column). Red boxes indicate recognized text areas, while blue arrows mark regions where text was missed. The number below each image represents the OCR metrics relative to the ground truth (GT) recognition result shown in the last column.
Figure S6: Qualitative evaluation on the DocUNet benchmark. Each panel shows the dewarped result with the corresponding horizontal AAD metric heatmap overlaid, allowing for a detailed comparison of correction errors among different methods.
Figure S7: Qualitative evaluation on the DIR300 benchmark. Each panel shows the dewarped result with the corresponding horizontal AAD metric heatmap overlaid, allowing for a detailed comparison of correction errors among different methods.
Figure S8: Qualitative evaluation on the UVDoc benchmark. Each panel shows the dewarped result with the corresponding horizontal AAD metric heatmap overlaid, allowing for a detailed comparison of correction errors among different methods.